OpenStudy (anonymous):

MEDAL! Matt plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Matt needs is $58.07. Which of the following is closest to the width of the portion of the driveway on which Matt plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 0.5 feet 1.5 feet 3 feet 6 feet

JoelTheBoss (joel_the_boss):

1 cubic yard = 27 cubit feet then we must find the no.of cubic yards .... for 1 cubic yard = $98 for "x" cubic yard = $58.07 1/x =98/58.07 x = 0.5925 Now we have to convert this in to cubic feet... 1 cubic yard = 27 cubic feet 0.5925 cubic yard = x cubic feet 1/0.5925= 27/x x = 15 .99 cubic feet now width x height x length = 16 cubic feet w x 8 x (1/3) = 16 w = 48/8 = 6 feet There for the width of the yard is approximately 6 feet!

OpenStudy (anonymous):

Wow, thank you so much! :)

JoelTheBoss (joel_the_boss):

your welcome :)